# InvBench: Can LLMs Accelerate Program Verification with Invariant Synthesis?

### v1

@@ Abstract
Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge. We introduce a principled framework for evaluating LLMs on invariant synthesis. Our approach uses a verifier-based decision procedure with a formal soundness guarantee and assesses not only correctness but also the speedup that invariants provide in verification. We evaluate 7 state-of-the-art LLMs, and existing LLM-based verifiers against the traditional solver UAutomizer. While LLM-based verifiers represent a promising direction, they do not yet offer a significant advantage over UAutomizer. Model capability also proves critical, as shown by sharp differences in speedups across models, and our benchmark remains an open challenge for current LLMs. Finally, we show that supervised fine-tuning and Best-of-N sampling can improve performance: fine-tuning on 3589 instances raises the percentage of speedup cases for Qwen3-Coder-480B from 8% to 29.2%, and Best-of-N sampling with N=16 improves Claude-sonnet-4 from 8.8% to 22.1%.
@@ 1 Introduction
Discovering such invariants is difficult and undecidable in general, which has motivated a long line of research. Traditional approaches include constraint solving (Colon et al., 2003; Gupta et al., 2009), dynamic analysis (Le et al., 2019), etc. Since invariant discovery is undecidable in general, researchers have tried a variety of learning-based methods (Li et al., 2017; Ezudheen et al., 2018). Building on this progression, the strong capabilities of LLMs in code generation and program reasoning (Austin et al., 2021; Chen et al., 2021; Wei et al., 2025b; c) naturally motivate a systematic evaluation of their potential for invariant discovery.
Pei et al. (2023) is the first work to evaluate the capabilities of LLMs in invariant generation. However, their methodology considers only correctness and does not assess how strong the generated invariants are. As a result, LLMs may generate correct invariants that perform well under their evaluation metric but provide no benefit for accelerating the verification process in real-world settings. Furthermore, their notion of correctness is not based on formal verification. Instead, it is determined by direct comparison with invariants generated by an existing tool, namely Daikon (Ernst et al., 2007). Daikon is a dynamic analysis tool whose invariants are not guaranteed to be sound, since they are inferred from observed test executions rather than proven across all possible executions. As a result, the ground-truth invariants themselves may be incorrect. Moreover, directly comparing LLM-generated invariants with Daikon’s output can lead to rejecting many correct invariants. For example, if an LLM proposes $a>0$ for an integer $a$ while Daikon reports $a\geq 1$, the evaluation in Pei et al. (2023) would incorrectly classify the LLM’s result as wrong, even though the two are equivalent. Hence, prior work’s evaluation methodology cannot reliably capture the correctness of LLM-generated invariants, let alone how useful they are for verification.
A series of follow-up works inspired by Pei et al. (2023) have proposed LLM-based verifiers (Wu et al., 2024b; a; Kamath et al., 2024). Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs. However, the evaluation of such tools suffers from two key limitations: (1) each work introduces a customized dataset and reports results only on that dataset, making cross-comparison impossible; and (2) many omit direct comparison with state-of-the-art traditional verifiers, underscoring the need for evaluation against established baselines.
@@
Figure 1: Illustration of InvBench’s evaluation pipeline. The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3). The verification procedure then incorporates this invariant to prove the property x != 145 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
Our work, InvBench, introduces a principled methodology for evaluating the capabilities of LLMs in invariant synthesis. Instead of checking whether LLMs reproduce invariants discovered by other tools, we employ a verifier-based decision procedure that directly determines the correctness of LLM-generated invariants, and we prove this procedure to be sound. We formalize the methodology as a proof calculus, providing a rigorous foundation for invariant evaluation. Unlike previous work that focuses on optimizing verification performance by designing tailored interactive protocols between solvers and LLMs (Chakraborty et al., 2023; Wu et al., 2024b; Kamath et al., 2024; Wu et al., 2024a), our goal is to develop a simple, first-order verification procedure suitable for evaluating the invariant generation capability of LLMs. Our formalization provides such a framework, offering a general solution for invariant evaluation.
Since the purpose of invariant synthesis is to accelerate verification, invariants that are too weak to aid verification or too difficult to verify as correct offer little practical value. To capture this, we evaluate the invariants by measuring the speedup they provide in the overall verification process.
To support comparison across solvers and LLMs, we construct a dataset of 226 instances derived from the most recent edition of the software verification competition SV-Comp (Beyer & Strejček, 2025) and use it to evaluate multiple LLM-based verifiers on this common benchmark. In addition, we assess the state-of-the-art traditional (i.e., non-LLM-based) verifier UAutomizer (Schüssele et al., 2024) both on our dataset and on each of the customized datasets introduced in prior work (Wu et al., 2024b; Kamath et al., 2024; Wu et al., 2024a).
UAutomizer consistently outperforms prior LLM-based verifiers on both our dataset and their customized datasets across all settings, with the only exception of LEMUR (Wu et al., 2024b), which was specifically designed for problems that UAutomizer fails to solve. These results suggest that while LLM-based verifiers are a promising direction, they do not yet offer a significant advantage over non-LLM-based approaches in general. Our findings further highlight the importance of using established datasets and baselines for meaningful progress in the community. They also highlight the role of the underlying solver, suggesting that base solver choice can influence outcomes.
In addition, we evaluate 7 state-of-the-art LLMs. The results indicate that model capability is critical, as shown by the sharp differences in achieved speedups. Nonetheless, LLMs remain largely unable to solve the hard subset of InvBench, underscoring that our dataset presents an open challenge for current LLMs.
We also investigate the extent to which supervised fine-tuning and Best-of-N sampling can enhance model performance in accelerating verification. To this end, we construct a fine-tuning dataset of 3589 instances and show that training on this dataset raises the percentage of speedup cases for Qwen3-Coder-480B from 8% to 29.2%. Similarly, Best-of-N sampling with N = 16 improves Claude-sonnet-4 from 8.8% to 22.1%.
@@
  • We propose a verification procedure for evaluating the invariant generation capabilities of LLMs, assessing both correctness and their effectiveness in accelerating verification.
  • We evaluate 7 state-of-the-art LLMs, and provide comparisons between existing LLM-based verifiers and the state-of-the-art non-LLM-based solver UAutomizer.
  • We construct a dataset of 3589 instances for training. We demonstrate that both supervised fine-tuning and Best-of-N sampling can improve model performance in accelerating verification.
@@ 2 Related Work
A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan & Qadeer, 2002; Lahiri & Bryant, 2007; Hojjat & Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot & Cousot, 1977; Cousot & Halbwachs, 1978; Cousot & Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala & McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich & Bodík, 2018). Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019). Daikon executes the program, observes runtime values, and reports properties that consistently hold over the observed executions. However, such invariants may fail to generalize to all possible executions, thereby compromising soundness. Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
Learning-Based Method for Invariant Generation
Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014; 2016; Ezudheen et al., 2018; Riley & Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020). More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a; b; d), giving rise to a series of work that explore using LLMs for finding invariants. Pei et al. (2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation. Chakraborty et al. (2023) proposes ranking the invariants generated by LLMs. LaM4Inv (Wu et al., 2024a) proposes a “query-filter-reassemble” strategy to leverage the generation power of LLMs, LEMUR (Wu et al., 2024b) proposes a sound algorithm with back-tracking. Loopy (Kamath et al., 2023) builds an LLM-based pipeline that combines clause sampling, Houdini filtering, and error-guided repair. Liu et al. (2024) proposes a new dataset including problems from courses and real-world systems. Our work explores fine-tuning and the best-of-N sampling method to improve LLMs’ capabilities.
@@ 3.1 Preliminary
We formalize the task of loop invariant synthesis using standard Hoare logic Hoare (1969). A Hoare triple $\{P\}\,S\,\{Q\}$ specifies that if the precondition $P$ holds before executing a statement $S$, then the postcondition $Q$ will hold after its execution. In the context of loops, an invariant $I$ is a logical proposition that summarizes the state of the program at each iteration, and it is the key to proving the validity of Hoare triples involving loops. For a loop of the form while $B$ do $S$, the goal of invariant synthesis is to identify a loop invariant $I$ that satisfies the following inference rule:
@@ 3.2 Verifier-Based Decision Procedure
$V(P,A,p)=\mathbf{T}\;\Rightarrow\;P\models_{A}p,\qquad V(P,A,p)=\mathbf{F}\;\Rightarrow\;P\not\models_{A}p.$
@@
$d_{a}:=V(P,\varnothing,q)\quad\text{(checking whether $q$ is a correct predicate),}$
[…]
An example of the two verifier queries are given in Figure˜1. The outcome of the procedure is expressed as a judgment
@@
The proof is provided in Section˜A.1. This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct. The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
@@ 3.3 Implementation
We describe the implementation of our verifier-based evaluation framework. Given a program $P$ and a target property $p^{\star}$, the system must generate candidate invariants $q$ and evaluate them according to the decision procedure. When proposing an invariant $q=\langle\psi,\ell\rangle$, the model selects a program location $\ell$ and supplies the corresponding predicate $\psi$.
@@
Parallel Evaluation. For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$. These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification. The final outcome is then derived exactly according to the decision calculus.
@@ 4 Experimental Setup
Dataset Split Avg. #Lines #Instances
Evaluation Easy 51 113
Hard 62 113
@@
Table 1: Dataset statistics of InvBench.
Dataset from SV-COMP.
We construct our benchmark from SV-COMP (Beyer & Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis. We collect a pool of 899 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record the solving time for each. Based on this runtime, we classify instances into an easy split (solved within 30 seconds) and a hard split (solved between 30 and 600 seconds). From this pool, we randomly sample 113 problems from each split, resulting in 226 instances in the final evaluation set. The selection process is fully automated with no manual cherry-picking.
Synthetic Dataset.
For LLM fine-tuning, we construct a synthetic corpus of verification problems using GPT-4o. Seed programs supplied in the prompt are selected to ensure no overlap with the 226 instances in the evaluation set, to avoid data leakage between training and evaluation. Each synthetic program is analyzed by UAutomizer, and we retain only those whose assertions are successfully proved. The invariants extracted from UAutomizer ’s execution logs constitute our final fine-tuning dataset, which contains 3589 problems paired with their invariants.
@@
All speedup-related measurements are reported relative to the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024), which serves as the baseline solver. We evaluate LLMs along two dimensions: the correctness of the generated invariants and the performance improvements they provide. Correctness is judged by the decision procedure formalized in Section˜3 with a timeout set to the problem’s original solving time by UAutomizer. For comparisons between UAutomizer and other tools, we report the number of solved instances under varying time budgets. We include the model’s token generation time in all evaluations.
@@
We benchmark Claude models from Anthropic, GPT models from OpenAI, and the Qwen family of models (Hui et al., 2024; Yang et al., 2025).
@@
Experiments were conducted on a server with Intel Xeon Platinum 8275CL CPUs (96 cores), 8 NVIDIA A100 GPUs, and 1.1 TB of memory, running Ubuntu 22.04.

### v2

# Quokka: Accelerating Program Verification with LLMs via Invariant Synthesis

@@ Abstract
Program verification relies on loop invariants, yet automatically discovering strong invariants remains a long-standing challenge. We investigate whether large language models (LLMs) can accelerate program verification by generating useful loop invariants. We introduce Quokka, a first-order and effective framework for LLM-based invariant synthesis that provides sound evaluation while achieving state-of-the-art speedup results. Unlike prior work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure. We construct a benchmark of 866 instances and evaluate 9 state-of-the-art LLMs across multiple model families. Our results show that Quokka consistently outperforms all prior LLM-based verifiers: achieving speedups of at least 1.2× on 81 instances compared to 39 instances for the previous best approach. We further demonstrate that supervised fine-tuning and Best-of-N sampling can yield measurable improvements in accelerating verification.
Machine Learning, ICML
@@ 1 Introduction
Discovering such invariants is difficult and undecidable in general, which has motivated a long line of research. Traditional approaches include constraint solving (Colon et al., 2003; Gupta et al., 2009), dynamic analysis (Le et al., 2019), etc. Since invariant discovery is challenging, researchers have tried a variety of learning-based methods (Li et al., 2017; Ezudheen et al., 2018). Building on this progression, the strong capabilities of LLMs in code generation and program reasoning (Austin et al., 2021; Chen et al., 2021; Wei et al., 2025b) naturally motivate a systematic evaluation of their potential for invariant discovery.
The first work (Pei et al., 2023) to evaluate LLMs for invariant generation has two key limitations. One limitation is that it only checks correctness without assessing whether invariants are strong enough to accelerate verification. Another limitation is that correctness is determined by directly comparing against Daikon (Ernst et al., 2007), a dynamic analysis tool that infers invariants from test executions rather than formal verification. This approach is unsound: Daikon’s invariants may themselves be incorrect. Moreover, semantically equivalent invariants (e.g., $a>0$ vs. $a\geq 1$ for integers) can be incorrectly rejected as wrong. Consequently, prior work cannot reliably evaluate either the correctness or the practical utility of LLM-generated invariants.
A series of follow-up works have proposed LLM-based verifiers. Instead of evaluating LLMs in isolation, these efforts develop verification frameworks powered by LLMs. This line of work predominantly adopts complex, sequential, and iterative processes with highly customized algorithms to handle LLM-generated invariants. For instance, LaM4Inv (Wu et al., 2024a) uses a “query-filter-reassemble” strategy that filters and combines model-generated predicates through complex logical operations (conjunctions or disjunctions) to construct valid loop invariants, rather than using them directly. LOOPY (Kamath et al., 2024) employs a customized Houdini algorithm with iterative refinement to filter candidate invariants. LEMUR (Wu et al., 2024b) develops a backtracking and sequential algorithm. While these approaches demonstrate the potential of LLM-based verification, their complexity raises a question: can we design a simpler, principled, and more effective decision procedure for LLM-based verifiers?
@@
Figure 1: Illustration of Quokka’s evaluation pipeline. The LLM proposes an invariant by specifying a program location and predicate (e.g., location B with x % 7 == 3). The verification procedure then incorporates this invariant to prove the property x != 700 using two verifier queries, and we measure the resulting speedup relative to a baseline without LLM assistance.
In this work, we introduce Quokka, a framework for LLM-based invariant synthesis that accelerates program verification. Quokka provides a sound approach to evaluate LLMs’ capabilities in generating invariants while achieving state-of-the-art speedup results compared with prior LLM-based verifiers. We formalize our methodology as a proof calculus, providing a rigorous foundation for building LLM-based verifiers. Instead of checking the correctness of invariants based on string-matching, we adopt a verifier-based approach that directly queries the underlying solver. Unlike previous work that designs complex, highly customized algorithms, Quokka employs a simple and principled verification procedure. Our approach enables both soundness guarantees and precise measurement of performance speedup, allowing us to systematically evaluate whether LLM-generated invariants can accelerate verification in practice.
To support systematic comparison across solvers and LLMs, we construct a dataset of 866 instances derived from the most recent edition of the software verification competition SV-COMP (Beyer and Strejček, 2025). To the best of our knowledge, this is the largest evaluation dataset for LLM-based verifiers to date, compared with datasets used in previous work. Quokka consistently outperforms all prior LLM-based verifiers. When comparing against LEMUR (the previous best LLM-based verifier with the same underlying solver), Quokka achieves speedups of at least 1.2× (i.e., verification completes at least 20% faster than the baseline solver) on 81 instances compared to LEMUR’s 39 instances. For more substantial speedups of at least 2.0× (i.e., verification completes at least twice as fast), Quokka reaches 51 instances compared to LEMUR’s 22 instances.
We evaluate 9 state-of-the-art LLMs, including GPT models from OpenAI, Claude models from Anthropic, and models from the Qwen and LLaMA families. Our experiments reveal that gpt-5.2 achieves the strongest speedup results among all evaluated models. Beyond model evaluation, we investigate techniques to enhance invariant generation capabilities. We construct a training dataset of 3589 instances using a verifier-based filtering approach that ensures data quality. Our experiments demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification. Specifically, Best-of-8 sampling achieves speedups of at least 1.2× on 99 instances, representing a 22% improvement over single-sample generation, which achieves such speedups on only 81 instances.
@@
  • We introduce a decision procedure for LLM-based invariant synthesis that provides sound evaluation and achieves the best speedup results compared with prior LLM-based verifiers.
  • We construct a benchmark of 866 instances, and conduct an evaluation of 9 state-of-the-art LLMs across multiple model families.
  • We construct training data and demonstrate that both supervised fine-tuning and Best-of-N sampling yield measurable improvements in accelerating verification.
@@ 2 Related Work
A long line of research has explored invariant synthesis using traditional techniques without machine learning, including model checking (Flanagan and Qadeer, 2002; Lahiri and Bryant, 2007; Hojjat and Rümmer, 2018; Vediramana Krishnan et al., 2024), abstract interpretation (Karr, 1976; Cousot and Cousot, 1977; Cousot and Halbwachs, 1978; Cousot and Cousot, 1979), constraint solving (Gulwani et al., 2009; Gupta et al., 2009), Craig interpolation (Jhala and McMillan, 2006; McMillan, 2010), and syntax-guided synthesis (Fedyukovich and Bodík, 2018). Prior work evaluating LLM-generated invariants (Pei et al., 2023) has relied on Daikon (Ernst et al., 2007), a tool for dynamic invariant detection (Echenim et al., 2019; Le et al., 2019). Daikon executes the program, observes runtime values, and reports properties that consistently hold over the observed executions. However, such invariants may fail to generalize to all possible executions, thereby compromising soundness. Our approach instead employs a verifier-based decision procedure relying on UAutomizer (Schüssele et al., 2024) that ensures soundness.
Learning-Based Method for Invariant Generation.
Machine learning based techniques have been widely adopted in invariant synthesis, including decision tree (Garg et al., 2014, 2016; Ezudheen et al., 2018; Riley and Fedyukovich, 2022; Xu et al., 2020), support vector machine (Li et al., 2017; Sharma et al., 2012), reinforcement learning (Si et al., 2018; Yu et al., 2023), and others (Sharma et al., 2013; Ryan et al., 2019; Yao et al., 2020). More recently, large language models have demonstrated strong capabilities in reasoning about code and logic (Wei et al., 2025a, b, c), giving rise to a series of work that explore using LLMs for finding invariants. (Pei et al., 2023) is the first pioneering work that evaluates LLMs’ capabilities in finding invariants, but it is not a sound evaluation. Various techniques have been proposed to couple LLMs with symbolic solvers, including ranking LLM-generated invariants (Chakraborty et al., 2023), the “query-filter-reassemble” strategy of LaM4Inv (Wu et al., 2024a), the back-tracking algorithm in LEMUR (Wu et al., 2024b), and Loopy’s integration of the classic Houdini algorithm (Kamath et al., 2024). On the dataset side, (Liu et al., 2024) introduces a rule-based method for constructing a fine-tuning corpus, which differs from our verifier-based approach. In contrast, our work provides a simple and sound evaluation procedure for assessing LLM-generated invariants and investigates how both fine-tuning and Best-of-N sampling can enhance LLM performance in invariant synthesis.
@@ 3.1 Preliminary
We formalize the task of loop invariant synthesis using standard Hoare logic (Hoare, 1969). A Hoare triple $\{P\}\,S\,\{Q\}$ specifies that if the precondition $P$ holds before executing a statement $S$, then the postcondition $Q$ will hold after its execution. In the context of loops, an invariant $I$ is a logical proposition that summarizes the state of the program at each iteration, and it is the key to proving the validity of Hoare triples involving loops. For a loop of the form while $B$ do $S$, the goal of invariant synthesis is to identify a loop invariant $I$ that satisfies the following inference rule:
@@ 3.2 Verifier-Based Decision Procedure
$V(P,A,p)=\mathbf{T}\;\Rightarrow\;P\models_{A}p,$
$V(P,A,p)=\mathbf{F}\;\Rightarrow\;P\not\models_{A}p.$
@@
$d_{a}:=V(P,\varnothing,q)$
$d_{b}:=V(P,\{q\},p^{\star})$
The query $d_{a}$ verifies whether $q$ is a correct invariant, while the query $d_{b}$ determines whether the target property holds under the assumption that $q$ is true. An illustration of these two verifier queries is provided in Figure˜1: $d_{a}$ corresponds to assert(x % 7 == 3) (upper subfigure), and $d_{b}$ corresponds to verifying the final assertion with the assume statement inserted (lower subfigure). The outcome of the procedure is expressed as a judgment
@@
Proof.
For the $\mathbf{T}$ case, the final rule must be DEC-PROP with $d=\mathbf{T}$. The premises yield $V(P,\varnothing,q)=\mathbf{T}$ and $V(P,\{q\},p^{\star})=\mathbf{T}$. By verifier soundness, $P\models q$ and $P\models_{\{q\}}p^{\star}$. Since $q$ holds on all executions of $P$, introducing the assumption $\{q\}$ does not remove executions relevant to $p^{\star}$; thus $P\models p^{\star}$. For the $\mathbf{F}$ case, the final rule must be DEC-FALSE. Its premise $V(P,\{q\},p^{\star})=\mathbf{F}$ implies $P\not\models_{\{q\}}p^{\star}$ by soundness. Assumptions restrict behaviors; hence, a violation under assumptions entails a violation without them, yielding $P\not\models p^{\star}$. ∎
This theorem establishes that whenever the calculus derives a conclusive outcome, that outcome is correct. The inconclusive case $\mathbf{U}$ is deliberately conservative: it makes no claim about the truth or falsity of the property and safely captures verifier incompleteness or timeouts.
@@ 3.3 Implementation
We describe the implementation of our verifier-based evaluation framework. Given a program $P$ and target property $p^{\star}$, the system evaluates the generated invariants according to the decision procedure. Each invariant $q=\langle\psi,\ell\rangle$ consists of a program location $\ell$ and predicate $\psi$.
@@
Parallel Verifier Queries. For each candidate $q$, the procedure issues two verifier queries, namely $d_{a}=V(P,\varnothing,q)$ to check whether $q$ is an invariant and $d_{b}=V(P,\{q\},p^{\star})$ to check whether the target holds under the assumption $q$. These queries are executed in parallel in our implementation, which reduces latency and enables speedup when the proposed invariant is useful for verification. The final outcome is then derived exactly according to the decision calculus.
3.4 Supervised Fine-Tuning and Best-of-N Sampling
Below, we discuss how we construct our dataset for fine-tuning and the way we perform Best-of-N sampling.
Synthetic Dataset Generation.
To construct the synthetic dataset, we prompt GPT-4o using the template in Appendix A.1. The template takes three seed programs as examples and instructs the model to synthesize a new C program that is compilable and contains both loops and assertions. To obtain a diverse and large pool of candidates, we repeatedly invoke the model with different seed programs. To avoid data leakage, these seed programs are randomly drawn from the SV-COMP pool  (Beyer and Strejček, 2025) that is disjoint from our evaluation set. Although the prompt requests compilable programs with loops and assertions, the LLM-generated programs may fail to compile, include assertions that do not hold, and contain multiple assertions. For programs with multiple assertions, we split them into separate instances, each retaining only a single assertion while preserving all loop structures (ensuring at least one loop per instance). We then run UAutomizer on every program and discard any instance that is non-compilable or whose assertion is invalid. This filtering step ensures the quality of the dataset, resulting in 3589 synthetic programs.
Extract Invariants Generated from UAutomizer.
When running UAutomizer to prove the assertions in the synthetic programs, the tool also emits the invariants it discovers. From its output, we extract the loop invariants. Each extracted invariant includes its program location and its predicate. Although each program contains exactly one assertion, it may include multiple loops, so a single program can yield multiple loop invariants, all associated with the same assertion. We pair each program with each of its corresponding loop invariants to form our training dataset. We show an example invariant generated from UAutomizer in Appendix A.2. With the ground-truth invariants, we perform supervised fine-tuning using LoRA (Hu et al., 2022).
Best-of-N Sampling.
Best-of-N sampling is an inference-time strategy in which multiple candidate programs are generated, and the most effective one is selected, a technique shown to improve performance on code-generation tasks (Ehrlich et al., 2025). In our setting, the best candidate is the invariant that yields the largest speedup, i.e., the one whose decision procedure finishes earliest. As described in Section˜3.2, evaluating a single candidate requires two verifier queries issued in parallel; therefore, Best-of-N sampling evaluates 2N verifier queries concurrently.
@@ 4 Experimental Setup
Dataset.
We construct our evaluation benchmark from SV-COMP (Beyer and Strejček, 2025), a standard competition in software verification, focusing on problems that require loop invariant synthesis. We collect a pool of 866 instances and run the state-of-the-art non-LLM verifier UAutomizer (Schüssele et al., 2024) with a 600-second timeout to record per-instance solving times. The process is fully automated, with no manual filtering or cherry-picking. We use this timeout to enable meaningful measurement of verification speedups. In contrast to prior work (Wu et al., 2024b), which evaluates only instances that the baseline cannot solve within this time budget, our approach requires knowing the baseline runtime to quantify relative improvements. Restricting to instances whose baseline runtime is observable allows us to report concrete speedup factors, rather than only counts of additionally solved problems. To the best of our knowledge, this dataset is the largest used in prior evaluations of LLM-based verifiers. Also, the programs in our dataset are significantly longer, with an average of 53 lines of code, compared to only 22 for LaM4Inv (Wu et al., 2024a), 23 for LEMUR (Wu et al., 2024b), and 27 for Loopy (Kamath et al., 2023). Manual inspection shows that our dataset contains features such as multiple loops, functions, arrays, and pointers, which are largely absent from prior datasets. This makes Quokka a more challenging and realistic benchmark that better distinguishes solver performance.
For training data, we construct our training data with synthetic data generation filtered by the solvers, as described in Section˜3.4. We make sure that the seed programs used for synthetic data generation are disjoint from our evaluation dataset, mitigating the leakage issue. The statistics of our dataset are shown in Table˜1.
Dataset Avg. #Lines #Instances
Evaluation 53 866
@@
Table 1: Dataset statistics.
Model # Correct Invariants # Fast1.2 # Fast1.4 # Fast1.6 # Fast1.8 # Fast2.0
Qwen2.5-7B 419 34 26 23 20 18
o3 550 38 33 30 24 23
Llama-3.1-8B 342 43 35 32 29 27
claude-opus-4.1 487 56 45 38 35 31
Qwen2.5-72B 500 63 49 45 37 33
claude-opus-4.5 689 69 53 46 39 34
gpt-5.1 694 56 46 41 40 37
claude-sonnet-4 620 68 58 51 45 41
gpt-5.2 710 81 65 60 56 51
Table 2: Performance of different LLMs when using Quokka. We report the number of instances with verified-correct invariants and the number of instances achieving speedups of at least $1.2\times$, $1.4\times$, $1.6\times$, $1.8\times$, and $2.0\times$ over UAutomizer.
@@
We evaluate LLMs along two dimensions: (1) the correctness of the generated invariants and (2) the performance improvements they provide. Correctness is determined by the solver using a timeout set to $1.2\times$ the original solving time for each problem. This slack allows the solver additional time to assess the correctness of model-generated invariants, reducing inconclusive outcomes caused by timeouts. All speedup-related measurements are reported relative to the original solver, which serves as the baseline. We report speedup using the metric of $\mathtt{\#Fast}_{p}$, defined as the number of instances for which verification completes at least $p\times$ faster than the baseline, with $p\in\{1.2,1.4,\ldots\}$. Additionally, we also report the number of instances solved under varying timeout budgets.
Choice of Solvers.
We use the state-of-the-art non-LLM-based solver UAutomizer (Schüssele et al., 2024) as our base solver. Our methodology requires that the solver can verify programs without externally-provided invariants, enabling us to measure speedup relative to a functional baseline. While prior work has explored other solvers such as Frama-C (Cuoq et al., 2012) and ESBMC (Gadelha et al., 2018), we select UAutomizer for two key reasons. First, UAutomizer represents the current state-of-the-art in automated verification, outperforming ESBMC on standard benchmarks in SV-COMP. Second, unlike Frama-C, which requires externally-provided invariants and cannot proceed without them, UAutomizer has built-in invariant synthesis capabilities that allow it to verify programs autonomously. To demonstrate the generality of Quokka’s methodology across different solvers, we also present experiments with ESBMC in Appendix A.5, where we show that Quokka successfully accelerates ESBMC as well.
@@
We benchmark Claude models from Anthropic, GPT models from OpenAI, and models from the Qwen and LLaMA families (Hui et al., 2024; Yang et al., 2025; Touvron et al., 2023).
@@
Experiments were conducted on a node with Intel E5-2640 v4 CPUs, 256G main memory, running Ubuntu 20.04 LTS.